Weatherly Lumber Company processes wood pulp for manufacturing various paper products. The company employs a process costing system for its manufacturing operations. All direct materials are added at the beginning of the process, and conversion costs are incurred uniformly throughout the process. This is the company’s production schedule for May:

Percent Completed

Tons of Pulp

Materials

Conversion

Work-in-Process Inventory, May 1

5,000

100

%

50

%

Started during May

16,000

Units to account for

21,000

Units from beginning Work-in-Process, which
were completed and transferred out during May

5,000

Started and completed during May

9,600

Work-in-Process Inventory, May 31

6,400

100

%

50

%

Total units accounted for

21,000

The following cost data are available:

Work-in-Process Inventory, May 1

Direct materials

$

79,350

Conversion

25,728

Costs incurred during May

Direct materials

253,920

Conversion

455,940

Required:

1. Calculate the equivalent units of direct materials and conversion during May. Use the FIFO method.

2. Calculate the cost per equivalent unit for both direct materials and conversion during May. Use the *FIFO* method.
Answer #1

1) Equivalent unit

Direct material Conversion Cost
Beginning WIP 5000*0% = 0 5000*50% = 2500
Unit started and completed 9600 9600
Ending WIP 6400 6400*50% = 3200
Equivalent unit 16000 15300

2) Cost per equivalent unit

Direct material = 253920/16000 = 15.87

Conversion = 455940/15300 = 29.80